# Bridging Voting and Deliberation with Algorithms: Field Insights from vTaiwan and Kultur Komitee

### v1

@@ 3.2.1. Motivation
The Human-in-the-Loop MES is a practical extension of the Method of Equal Shares algorithm, co-designed with KK24 for their needs, to enhance transparency and build algorithmic trust by incorporating real-time digital feedback. This method allows participants to control the extent to which decision-making is delegated to the voting aggregation algorithm versus reserved for deliberation.
3.2.2. Method of Equal Shares (MES)
The Greedy method, commonly used in standard participatory budgeting, selects projects with the highest votes but ignores proportionality. While simple and efficient, it favours majority preferences and overlooks minority groups. The Method of Equal Shares (MES), proposed by Peters et al. (2021), ensures proportional and fair resource allocation in collective decisions. Each voter is assigned an equal starting budget , determined iteratively. Here, denotes the utility voter derives from project , representing the perceived benefit or preference for the project. A project is -affordable for if its cost can be covered by voters contributing up to monetary units per unit of utility: MES proceeds as follows:
Check Affordability: Identify if any project is -affordable for some . A project is -affordable when its total cost can be funded by voters, each contributing no more than monetary units per unit of utility. If no project is affordable, the process ends, returning the set of winners .
Select and Deduct: Add the most affordable project to the set of winners , and update each voter’s budget:
This ensures that voters contribute only as much as their remaining budget allows, with contributions capped by their proportional share based on .
The starting budget begins at , where is the total budget and is the number of voters. Incremental adjustments to the budget allow flexibility to explore alternative project selections without exceeding the total budget . By iteratively selecting projects that align contributions with utility and budget constraints, MES guarantees proportionality. This ensures that groups with shared preferences are fairly represented in the selected set of projects.
@@ 3.3.2. Online Phase: Wiki-Survey with Polis
The Online Phase uses Polis, a participatory wiki-survey platform, to enable scalable deliberation. Participants can submit statements and vote by agreeing, disagreeing, or abstaining, fostering the identification of consensus and divisive opinions. As described in The Computational Democracy Project webpage 222https://compdemocracy.org/algorithms/ and work by Small et al. (2021), Polis uses clustering algorithms to analyse responses and generate real-time reports. Fine-grained clustering applies K-means at for real-time visualisation, while coarse-grained clustering iterates K-means for to , optimised using silhouette coefficients to ensure stability. The final report highlights group consensus and divisive statements, such as the identification of five opinion groups in the AI regulation deliberation in the vTaiwan Polis report333https://polis.tw/report/r3dvith8ntmwywyf4nctc.
@@ 3.3.3. Offline Phase: Decision Tree and Face-to-Face Deliberation
Building on the insights from the Online Phase, the proposed Offline Phase integrates digital data into structured physical deliberations. This phase, co-designed with vTaiwan, ensures that discussions are grounded in evidence while allowing participants to learn and refine their opinions through direct interaction.
Decision Tree Creation: The clusters and divisive statements identified in the Polis report are used to construct a decision tree. This tree visually represents how the crowd is divided by key issues, serving as a roadmap for discussions.
Deliberation Event: Structured deliberations based on divisive statements unfold in the following steps:
Pre-Discussion Voting: Participants use Mentimeter to vote on divisive statements via a 5-point Likert scale (e.g., from strongly disagree to strongly agree).
Discussion Round: Moderators facilitate dialogue by inviting participants from diverse opinion groups to share their perspectives and reasoning. This encourages open dialogue and mutual understanding.
Post-Discussion Voting: Participants re-vote on the same statements after discussions. Moderators prompt participants to reflect on and share reasons for any changes in their stance.
In this iteration, 104 participants voted online using Polis by 17th December 2024, while 44 self-invited participants attended a physical deliberation workshop on 20th December in Taipei. Although the two groups overlap, they are not subsets of each other, as some workshop attendees did not participate in the online voting. The two phases were conducted independently, with the online results serving to collect public opinion and guide the physical deliberation. No demographic data was collected.
@@ 3.3.4. Evaluation Metrics
To evaluate the impact of deliberation, we used four key metrics. The Percentage of Voters Changed captures the proportion of participants who shifted their stance. The Polarisation Index, measuring opinion dispersion, is calculated as the standard deviation of responses divided by the response range. The Consensus Index quantifies group alignment as the proportion of responses matching the majority opinion. Lastly, the Mean Opinion Change evaluates shifts in sentiment as the difference between the average post-deliberation and pre-deliberation scores. These metrics assess how deliberation influences opinions, reduces polarisation, and fosters consensus.
@@ 4. Data and Implementation
We apply a data-driven approach to quantitative voting data to understand voting patterns and preferences. This provides insights into the efficacy of these participatory mechanisms. Complementary, we evaluate qualitative feedback that was collected through interviews and surveys by the organising committees. This offers a deeper understanding of how these voting patterns influence deliberative outcomes, adding an additional layer of context to the quantitative data. The responses in German (KK24) and Traditional Chinese (vTaiwan) were translated using DeepL before the analysis.
As researchers advising citizen groups in this participatory research, we ensured academic integrity by clearly defining our roles as advisors and observers. In both roles, the processes were co-designed with the organising committees through multiple design meetings from late 2023 to 2024, which shaped the methods and tools to suit their needs. We guided the application of methods, built interfaces and explanations, and facilitated technical implementation while the organising committees and participants then took their decisions autonomously. The data analysed in this study are all secondary and anonymised, independently collected by the organising committees of KK24 and vTaiwan during their processes. No primary data collection was conducted by the research team. For KK24, the voting data and survey responses were gathered with signed consent after participants were briefed on the purpose, data use, privacy, and withdrawal rights. Data collected by vTaiwan using Polis and Mentimeter included no personal information, was also obtained with participant consent, and is publicly available as open data on vTaiwan’s GitHub444github.com/v-taiwan/241220-AI-Regulation.
@@ 5. Results
5.1. Radial Clustering for Preference Based Subgroups
@@ 5.1.1. Heterogeneous deliberation outcome mirror the voting outcome more closely
The p-values in Figure 5 reveal a significant correlation between individual voting outcomes and group-assigned deliberation points in the heterogeneous round (), but a non-significant correlation in the homogeneous round (). This suggests that in heterogeneous deliberation groups, where diverse perspectives are expected to emerge, the final group decisions closely mirror individual online voting outcomes. We also observed more groups using sticky notes as votes to make collective decisions in the heterogeneous round, as shown in Figure 7. This pattern is further illustrated in Appendix Table LABEL:tab:final_projects. Among the 17 projects selected during the deliberation process, 7 overlap with those that would have been chosen if online voting alone determined the funding allocation for this portion of the budget. Compared to a hypothetical pure voting scenario, heterogeneous deliberation introduces fewer unique projects. Only 3 projects were selected primarily due to heterogeneous deliberation (tagged HT), while 7 projects were selected primarily due to homogeneous deliberation (tagged HM). This suggests that homogeneous deliberation incorporates a wider range of projects that diverge from individual voting outcomes.
One possible explanation is that participants with diverse preferences in heterogeneous groups find it challenging to reach consensus, often resorting to in-group voting to resolve differences. These findings highlight the need to critically assess the impact of group dynamics in deliberative processes within systems where voting already aggregates preferences effectively.
@@ 5.1.2. Heterogeneous deliberation tends to fund more costly projects
As detailed in Table 13 in the appendix, the hypothetical outcomes represent the project selections if the 190,000 CHF budget were allocated to projects based solely on decisions from homogeneous or heterogeneous groups. Heterogeneous deliberation funded fewer but more expensive projects than homogeneous deliberation, as evidenced by higher mean (18,990 CHF vs. 13,350 CHF) and median (13,000 CHF vs. 11,450 CHF) project costs. These findings suggest that diverse perspectives may lead to prioritising larger or higher-impact projects.
@@
As shown in Figure 13, participant responses highlight notable differences in the perceived ease of decision-making and the extent to which outcomes reflected their preferences across the two rounds of deliberation. The first round, organised using Radial clustering to create homogeneous groups, was perceived as easier, with 83% of participants agreeing that reaching a decision was easy (52% “Somewhat agree” and 31% “Strongly agree”). 65% of the participants agreed in the second round, where Radial clustering was used to form heterogeneous groups (31% “Somewhat agree”, and 34% “Strongly agree”). Similarly, decisions made by the homogeneous groups were more likely to reflect participants’ preferences, with 76% agreeing (62% “Somewhat agree” and 14% “Strongly agree”), compared to 55% in the heterogeneous groups (34% “Somewhat agree” and 21% “Strongly agree”). These results suggest that using Radial clustering to organise groupings creates a perceivable difference for participants, with homogeneous groupings leading to smoother decision-making and outcomes more aligned with individual preferences.
@@ 5.1.4. Younger participants exhibit significant alignment differences between deliberation rounds
Figure 8 compares alignment, the proportion across homogeneous and heterogeneous deliberation rounds, segmented by gender and age groups. In the homogeneous round, younger participants (Age ) and female participants exhibit higher median alignment (0.33 and 0.35, respectively), while older participants (Age 42) and male participants have lower alignment medians (0.23 and 0.32, respectively). In the heterogeneous round, alignment for younger participants and females decreases (to 0.28 and 0.30), whereas older participants and males experience increases (to 0.34 and 0.37, respectively).
Due to the smaller sample sizes within each group, only the difference in alignment for younger participants meets the threshold for statistical significance (). For the other groups, the observed differences do not reach statistical significance ( for males, for females, for middle-aged participants, and for older participants).
@@ 5.1.5. The perceptions of group dynamics of deliberation were mixed
While most participants did not specifically comment on differences between the two rounds of deliberation, they perceived the deliberations differently. Some participants noted difficulties in ensuring equal participation within small groups. P01 observed, “When working in groups, the stronger voices sometimes had more say. I then tried to mediate, which was sometimes difficult.” P07 echoed these challenges but appreciated the structure provided by facilitators, noting, “The main challenge to contribute my opinion that I have developed or to stand up for it / to share myself in small groups… However, (KK24) they have made this as simple as possible with the structure.” Others highlighted the inclusivity fostered by group dynamics. P04 remarked positively, “Through the different groups, there was always the opportunity to share one’s voice.” Similarly, P21 appreciated the multiple avenues for participation, stating, “I find my opportunity to participate in decision-making to be appropriate overall since various opportunities for participation (voting, polls, meetings in smaller groups) were provided.” Participants also reflected on the balance between individual preferences and collective outcomes. P16 praised the alignment between decision-making and group interests, stating, “The decision-making process was good and based on the group’s own interests.” P27 emphasised the effective integration of individual and group perspectives, describing it as a “Very good mix of individual opinions and discussion.” While homogeneous groups may foster alignment by reducing conflicting perspectives, heterogeneous groups encourage engagement with diverse viewpoints, which may lead to some of these mixed feelings regarding the group dynamic. Future efforts should address the influence of dominant voices while enhancing structures that support fair participation and constructive dialogue.
@@ 5.2. Human-in-the-loop MES
5.2.1. MES distributes budget more fairly and fund more projects
In Figure 9, under the same budget constraint of 190,000 CHF, the Method of Equal Shares (MES) funds more affordable projects than the conventional and commonly used Greedy Method (taking projects with the highest votes), leading to a higher number of projects won per voter. While the budget allocation per voter is similar in both methods, MES distributes resources more fairly, giving more to voters who receive less under the Greedy Method. This redistribution lowers the Gini coefficient from 0.17 to 0.14, reflecting a fairer outcome. The light green lines highlight these changes, showing how MES benefits voters disadvantaged by the Greedy Method while staying within the same budget.
@@ 5.2.2. Participants value Human-in-the-loop MES for balancing algorithmic decisions and human input
After being shown different MES budget scenarios, participants opted to allocate 50% of the budget using MES. The interactive interface displayed 18 selected projects, and participants were divided into randomly assigned groups to discuss potential vetoes or budget adjustments. None of the groups vetoed any projects, indicating approval of the algorithmically selected set. However, participants proposed reducing the budget for the project “Kultur-Agenda,” an online calendar for art and cultural events, from 18,000 CHF to 9,000 CHF. This additional step allowed participants to refine algorithmic decisions, balancing automation with human judgment.
Figures 16 and 16 in the appendix present participant responses to two key survey questions about MES fairness and the balance between voting and deliberation. The responses indicate that participants generally perceived MES as fair, with 62% rating it “Very fair” and 23% rating it “Somewhat fair.” A majority expressed support for maintaining the 50:50 ratio of individual voting to group deliberation, with an overwhelming 81% preferring the same 50:50 proportion and only 15% and 3% advocating for having a higher portion of voting or deliberation, respectively.
In the open-ended feedback questions, participants generally expressed appreciation for the fairness and structure of MES but highlighted areas for improvement in preparation and communication. Many valued its ability to ensure all votes are counted equally, as P02 noted, “It is pragmatic and valuable to count every vote that is cast. Allows less verbal people to have equal weight given to their voice.” Others praised its practical approach to budget management, with P09 stating, “The Method of Equal Shares provided me with a reassuring guide.” P19 shared, “It’s nice that half of the budget is distributed automatically, and the rest is then discussed and decided in detail.” P03 commented on the balance of methods, saying, “The good thing was the mix of analogue and digital.” P27 added, “The best thing was the mixture of individual decisions and group decisions.” P21 acknowledged the trade-offs inherent in collective decision-making, commenting, “Since it was a large group, the opinions of individuals may have been lost a little, as many things were decided by voting, but I still felt heard and taken seriously.” Suggestions for enhancement included utilising online voting more extensively (P20: “More online voting; With this voting method (MES), there is no external influence.”).
While participants’ feedback did not focus much on the flexibility of Human-in-the-loop MES, since the 50:50 voting-deliberation ratio was established early on in the discussion, they frequently highlighted the value of combining voting with deliberation. This mix was seen as an effective way to balance individual input with collective decision-making, ensuring fairness and practicality.

### v2

@@ 3.2.1. Motivation
The Human-in-the-Loop Method of Equal Shares is a practical extension of the Method of Equal Shares algorithm, co-designed with KK24 for their needs, to enhance transparency and build algorithmic trust by incorporating real-time digital feedback. This method allows participants to control the extent to which decision-making is delegated to the voting aggregation algorithm versus reserved for deliberation.
3.2.2. Method of Equal Shares (MES)(Peters et al., 2021)
The Greedy method commonly used in participatory budgeting (PB) selects projects with the highest votes but ignores proportionality, favouring majority preferences while overlooking minority groups. Research by Yang et al. (2024a) shows participants in PB settings are rarely cost-conscious, leading to projects winning disproportionate budget shares relative to their votes. The Method of Equal Shares (MES) (Peters et al., 2021) ensures fair resource allocation by assigning each voter an equal starting budget . Here, represents the utility voter derives from project . MES operates through two key steps: (1) Check Affordability. A project is -affordable if its cost can be covered by voters contributing proportionally to their utility, with each contribution capped at . (2) Select and Update. The most affordable project is added to the winner set , and each voter’s remaining budget is updated accordingly. The algorithm begins with budget shares of (total budget divided by number of voters) and iteratively selects projects until no more are affordable. This approach guarantees proportionality, ensuring groups with shared preferences receive fair representation in the final project selection.
@@ 3.2.3. The Human-in-the-Loop extension
3.2.4. Evaluation Metrics
To evaluate Human-in-the-Loop MES, we employed metrics addressing both algorithmic fairness and human agency: project representation measured the number of projects won per voter across methods; budget allocation fairness was quantified using the Gini coefficient (), with lower values indicating more equitable distribution; perceived fairness was assessed through participant survey responses on a 5 point Likert scale; and human agency was measured by examining both the percentage of algorithmic decisions modified through deliberation and participant preferences for the voting-to-deliberation ratio.
@@ 3.3.2. Online Phase: Wiki-Survey with Polis
The Online Phase uses Polis, a participatory wiki-survey platform, to enable scalable deliberation. Participants can submit statements and vote by agreeing, disagreeing, or abstaining, fostering the identification of consensus and divisive opinions. As described in The Computational Democracy Project webpage 333https://compdemocracy.org/algorithms/ and work by Small et al. (2021), Polis uses clustering algorithms to analyse responses and generate real-time reports. These reports can include a low-dimensional visualization of the collected data and several summary statistics of the debate. For example, the final report in the AI regulation deliberation in vTaiwan444https://polis.tw/report/r3dvith8ntmwywyf4nctc identifies 5 opinion groups, and highlights 2 divisive and 14 consensus statements.
@@ 3.3.3. Offline Phase: Decision Tree and Face-to-Face Deliberation
Building on the online phase results, our offline phase integrated digital insights into structured physical deliberations. Co-designed with vTaiwan, this approach ensured evidence-based discussions while facilitating direct learning and opinion refinement.
Decision Tree Creation: We used clusters and divisive statements from the Polis report to construct a visual decision tree (Figure 12), mapping how participants divided on key issues. This served as a discussion roadmap, making complex opinion patterns accessible to moderators and participants.
Deliberation Process: The structured event followed three steps: First, participants used Mentimeter for pre-discussion voting on divisive statements using a 5-point Likert scale. Second, moderators facilitated dialogue by inviting participants from diverse opinion groups to share their perspectives, encouraging mutual understanding. Finally, participants re-voted on the same statements and reflected on reasons for any changes in their positions.
In this iteration, 104 participants contributed to the online Polis survey by 17th December, while 44 attended the physical workshop on 20th December in Taipei. The groups partially overlapped, with some workshop attendees not participating in the online phase. The phases operated independently, with online results informing but not constraining the physical deliberation agenda. No demographic data was collected.
@@ 3.3.4. Evaluation Metrics
To evaluate deliberation’s impact, we employed four key metrics: percentage of voters changed, measured as the percentage of voters who changed their stance after deliberation; polarisation reduction, quantified using the Bimodality Coefficient (BC = ), where values exceeding 0.555 indicate opinion polarisation (Knapp, 2007); consensus building, calculated as , with higher values indicating greater agreement; and mean opinion changes, measuring directional shifts in average responses on the 5-point Likert scale (-2 to +2). Statistical significance was assessed using paired t-tests for mean changes.
@@ 4. Data and Implementation
We employed mixed methods to analyse participatory decision-making processes, combining quantitative voting data with qualitative feedback. This approach provided insights into voting patterns, algorithmic performance, and participant experiences. German and Traditional Chinese responses were translated using DeepL to facilitate analysis.
As researchers engaging with these democratic innovations, we adopted distinct roles in each case study. For KK24, we served as academic advisors, co-designing the process with the organising committee through multiple meetings from late 2023 to 2024. We provided methodological guidance, built interfaces and offered technical implementation support while ensuring decision-making autonomy remained with the organisers and participants. In the vTaiwan case, the first author participated in the public deliberation workshops as a regular attendee while documenting the process, using only data that was already publicly available online. This approach allowed for observation of deliberative dynamics firsthand while maintaining analytical distance.
All data analysed in this study are secondary, anonymised, and collected by the respective organising committees. KK24 data were gathered with signed participant consent following comprehensive briefings on data use and privacy rights. vTaiwan data contained no personal information and were collected with participant consent through Polis and Mentimeter platforms. Both datasets are publicly available for research transparency and replication555KK24 dataset: github.com/joshuay1/kk24-deliberation-voting-dataset666vTaiwan dataset: github.com/v-taiwan/241220-AI-Regulation.
@@ 5. Results
5.1. Preference-based Clustering for Deliberation (PCD)
@@ 5.1.1. Heterogeneous deliberation outcome mirror the voting outcome more closely
The p-values in Figure 2 reveal a significant correlation between individual voting outcomes and group-assigned deliberation points in the heterogeneous round (), but a non-significant correlation in the homogeneous round (). This suggests that in heterogeneous deliberation groups, where diverse perspectives are expected to emerge, the final group decisions closely mirror individual online voting outcomes. We also observed more groups using sticky notes as votes to make collective decisions in the heterogeneous round, as shown in Figure 5.
This pattern is further illustrated in Table LABEL:tab:final_projects in the appendix. Among the 17 projects selected during the deliberation process, 7 overlap with those that would have been chosen if online voting alone determined the funding allocation for this portion of the budget. Compared to pure voting, heterogeneous deliberation introduces fewer unique projects. Only 3 projects were selected primarily due to heterogeneous deliberation (tagged HT), while 7 projects were selected primarily due to homogeneous deliberation (tagged HM). This suggests that homogeneous deliberation incorporates a wider range of projects that diverge from individual voting outcomes. One possible explanation is that participants with diverse preferences in heterogeneous groups find it challenging to reach consensus, often resorting to in-group voting to resolve differences. These findings motivate the study of group dynamics in deliberative processes within systems where voting already aggregates preferences effectively.
@@ 5.1.2. Heterogeneous deliberation tends to fund more costly projects
As detailed in Table 1, the hypothetical outcomes represent the project selections if the 190,000 CHF budget were allocated to projects based solely on decisions from homogeneous or heterogeneous groups. Heterogeneous deliberation funded fewer but more expensive projects than homogeneous deliberation. These findings suggest that diverse perspectives may lead to prioritising larger or higher-impact projects.
| Metric | Homogeneous | Heterogeneous |
| Total Budget | 186,900 | 189,900 |
| Number of Projects | 14 | 10 |
| Mean Cost | 13,350.00 | 18,990.00 |
| Cost Std. Dev. | 8,581.71 | 12,657.05 |
@@
As shown in Figure 6, participant responses highlight notable differences in the perceived ease of decision-making and the extent to which outcomes reflected their preferences across the two rounds of deliberation. The first round, organised using Radial clustering to create homogeneous groups, was perceived as easier, with 83% of participants agreeing that reaching a decision was easy. 65% of the participants agreed in the second round, where Radial clustering was used to form heterogeneous groups. Similarly, decisions made by the homogeneous groups were more likely to reflect participants’ preferences, with 76% agreeing, compared to 55% in the heterogeneous groups. These results suggest that using Radial clustering to organise groupings creates a perceivable difference for participants, with homogeneous groupings leading to smoother decision-making and outcomes more aligned with individual preferences.
@@ 5.1.4. Younger participants exhibit significant alignment differences between deliberation rounds
Figure 3 compares alignment, the proportion across homogeneous and heterogeneous deliberation rounds, segmented by gender and age groups. In the homogeneous round, younger participants (Age ) and female participants exhibit higher median alignment (0.33 and 0.35, respectively), while older participants (Age 42) and male participants have lower alignment medians (0.23 and 0.32, respectively). In the heterogeneous round, alignment for younger participants and females decreases (to 0.28 and 0.30), whereas older participants and males experience increases (to 0.34 and 0.37, respectively). Due to the smaller sample sizes within each group, only the difference in alignment for younger participants meets the threshold for statistical significance (). For the other groups, the observed differences do not reach statistical significance.
@@ 5.1.5. The perceptions of group dynamics of deliberation were mixed
While most participants did not specifically comment on differences between the two rounds of deliberation, they perceived the deliberations differently. Some participants noted difficulties in ensuring equal participation within small groups. P01 observed, “When working in groups, the stronger voices sometimes had more say. I then tried to mediate, which was sometimes difficult.” P07 echoed these challenges but appreciated the structure provided by facilitators, noting, “The main challenge to contribute my opinion that I have developed or to stand up for it / to share myself in small groups… However, (KK24) they have made this as simple as possible with the structure.” Others highlighted the inclusivity fostered by group dynamics. P04 remarked positively, “Through the different groups, there was always the opportunity to share one’s voice.” Similarly, P21 appreciated the multiple avenues for participation, stating, “I find my opportunity to participate in decision-making to be appropriate overall since various opportunities for participation (voting, polls, meetings in smaller groups) were provided.” Participants also reflected on the balance between individual preferences and collective outcomes. P16 praised the alignment between decision-making and group interests, stating, “The decision-making process was good and based on the group’s own interests.” P27 emphasised the effective integration of individual and group perspectives, describing it as a “Very good mix of individual opinions and discussion.” While homogeneous groups may foster alignment by reducing conflicting perspectives, heterogeneous groups encourage engagement with diverse viewpoints, which may lead to mixed feelings regarding the group dynamic. Future efforts should address the influence of dominant voices while enhancing structures that support fair participation and constructive dialogue.
@@ 5.2. Human-in-the-loop MES
5.2.1. MES distributes budget more fairly and funds more projects
Shown in Figure 8 in the appendix, under the same budget constraint of 190,000 CHF, the Method of Equal Shares (MES) funds more affordable projects than the conventional and commonly used Greedy Method (taking projects with the highest votes), leading to a higher mean number of projects won per voter, from 8.65 to 11.76. While the budget allocation per voter is similar in both methods, MES distributes resources more fairly, giving more to voters who receive less under the Greedy Method. This redistribution lowers the Gini coefficient from 0.17 to 0.14, reflecting a fairer outcome. The light green lines highlight these changes, showing how MES benefits voters disadvantaged by the Greedy Method while staying within the same budget.
| Statement | Voters Changed (%) | Before Deliberation | After Deliberation | Change | ||||||
| Mean | BC | CI | Mean | BC | CI | Mean | BC | CI | ||
| 1. Ban AI political propaganda | 36.7 | 0.00 | 0.622 | 0.41 | -0.23 | 0.543 | 0.42 | -0.23 | -0.079 | 0.01 |
| 2. AI companies should disclose | 53.1 | -0.03 | 0.561 | 0.43 | -0.34 | 0.344 | 0.50 | -0.31 | -0.217 | 0.07 |
| 3. AI’s industrial impact | 30.0 | -0.53 | 0.524 | 0.46 | -0.57 | 0.513 | 0.47 | -0.03 | -0.011 | 0.01 |
| 4. Governments should prevent bias | 37.9 | -0.17 | 0.660 | 0.41 | 0.69 | 0.619 | 0.44 | 0.86∗ | -0.040 | 0.02 |
| 5. AI developers must be accountable | 45.2 | -0.29 | 0.526 | 0.44 | -0.03 | 0.443 | 0.47 | 0.26 | -0.083 | 0.03 |
| Note: BC = Bimodality Coefficient (values ¿ 0.555 indicate bimodal distribution); CI = Consensus Index (1/(1+standard deviation)), higher values indicate more consensus; ∗ indicates statistically significant change (p ¡ 0.05). | ||||||||||
@@ 5.2.2. Participants value Human-in-the-loop MES for balancing algorithmic decisions and human input
After being shown different MES budget scenarios, participants opted to allocate 50% of the budget using MES. The interactive interface displayed 18 selected projects, and participants were divided into randomly assigned groups to discuss potential vetoes or budget adjustments. None of the groups vetoed any projects, indicating approval of the algorithmically selected set. However, in the budget adjustment session shown in Figure 5, participants proposed reducing the budget for the project “Kultur-Agenda,” an online calendar for art and cultural events, from 18,000 CHF to 9,000 CHF. This additional step allowed participants to refine algorithmic decisions, balancing automation with human judgment.
Figures 10 and 11 in the appendix present participant responses to two key survey questions about MES fairness and the balance between voting and deliberation. The responses indicate that participants generally perceived MES as fair, with 62% rating it “Very fair” and 23% rating it “Somewhat fair.” A majority expressed support for maintaining the 50:50 ratio of individual voting to group deliberation, with an overwhelming 81% preferring the same 50:50 proportion and only 15% and 3% advocating for having a higher portion of voting or deliberation, respectively.
In the open-ended feedback questions, participants expressed appreciation for the fairness and structure of MES but highlighted areas for improvement in preparation and communication. Many valued its ability to ensure all votes are counted equally, as P02 noted, “It is pragmatic and valuable to count every vote that is cast. Allows less verbal people to have equal weight given to their voice.” Others praised its practical approach to budget management, with P09 stating, “The Method of Equal Shares provided me with a reassuring guide.” P19 shared, “It’s nice that half of the budget is distributed automatically, and the rest is then discussed and decided in detail.” P03 commented on the balance of methods, saying, “The good thing was the mix of analogue and digital.” P27 added, “The best thing was the mixture of individual decisions and group decisions.” P20 suggested using more voting, stating: “More online voting; With this voting method (MES), there is no external influence.”). While participants’ feedback did not focus much on the flexibility of Human-in-the-loop MES, since the 50:50 voting-deliberation ratio was established early on in the discussion, they frequently highlighted the value of combining voting with deliberation.